OpenStudy (diamondanon2):

Solve F(x) for the given domain. Include all of your work in your final answer. Submit your solution. Solve F(x) for the given domain. Include all of your work in your final answer. Submit your solution. F(x) = x^2 + 2 F(x^2) =

OpenStudy (loser66):

\(f(\color{red}{x})=\color{red}{x}^2 +2\) Now \(f(\color{blue}{x^2})= ...+2\) Can you fill it up?

OpenStudy (diamondanon2):

2x^2 ? @Loser66

OpenStudy (diamondanon2):

What would the steps be to find that out?

OpenStudy (loser66):

ok, go steps!! if f(x) = x , then f(y) =?

OpenStudy (diamondanon2):

y

OpenStudy (diamondanon2):

@Loser66

OpenStudy (loser66):

yeah, now f(z) =?

OpenStudy (diamondanon2):

z

OpenStudy (loser66):

ok, a litlle bit harder if f(x) =x+2, then f(y) =?

OpenStudy (diamondanon2):

y = y + 2

OpenStudy (loser66):

f(y) = y+2, not y = y+2, ok?

OpenStudy (diamondanon2):

Is is x^2 + 4?

OpenStudy (diamondanon2):

and okay

OpenStudy (loser66):

so, if I say, if f(x) = x^2 +2, then , f(y) = ?

OpenStudy (diamondanon2):

f(y) = y^2 + 2

OpenStudy (loser66):

yup yup yup

OpenStudy (loser66):